# Demystifying Quantum Power Flow: Unveiling the Limits of Practical Quantum Advantage

### v2

@@ III Classical Method Complexity
The Conjugate Gradient (CG) method is among the best algorithms in terms of asymptotic complexity for solving linear systems of equations with a sparse, positive definite matrix $A$. As the power system matrices fulfill both conditions, we focus on analyze CG as one of the best ways to solve DCPF classically. We define $s$ as the maximum number of non-zero values in a row i.e. sparsity, and $\kappa$ as condition number ($\kappa=\|A^{-1}\|\|A\|$). The total complexity of CG can be split into two parts: complexity of dominant operations and number of times dominant operations are used.
Now, considering the dominant operation complexity of matrix-vector multiplication, $\mathcal{O}(Ns),$ and upper bound on number of iterations required to achieve the error $\varepsilon_{c}$ (details are given in supplementary information), the overall complexity of the CG for solving (1) is $\mathcal{O}(N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}}))$. Further, read-write complexity for the sparse matrix $A$ is $\mathcal{O}(Ns)$ (higher than $\mathbf{b},\mathbf{x}$ read-out costs which is $\mathcal{O}(N)$) and this additional cost is additive to the runtime complexity. Here, additive means that we read $A,\mathbf{b}$ once before solving (1) and read $\mathbf{x}\in\mathbb{R}^{N}$ once after solving the problem. To obtain $\mathbf{x}$ one does not need to repeat CG but only needs to read it with complexity $\mathcal{O}(N)$. This is important as additive terms are subsumed into dominant complexity terms and do not alter the end-to-end complexity. Thus, the end-to-end complexity of solving DCPF (1) using CG is $\mathcal{O}(Ns+N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}})+N)=\mathcal{O}(N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}}))$. [^4] For basic Big-O notation introduction, readers can refer these notes: Link. Further, along with exploiting sparsity, CG runtime complexity has square root dependence condition number $\kappa$, making it highly efficient for DCPF as $\kappa$ has polynomial growth from Figure 3 (center).
From the definition of $\varepsilon_{c}$ and error vectors ($\mathbf{e}^{o}$ and $\mathbf{e}^{i}$, details are given in supplementary material), it is clear that CG error upon convergence will depend on initial point $\mathbf{x}_{o}$. It is clear that if a starting point is closer to the final solution, the same value of $\varepsilon_{c}$ can be achieved in fewer CG iterations. This feature of CG has led to considerable efforts on finding ‘good’ starting points of phase-angles ($\mathbf{x}$) in DCPF problems, particularly when DCPF needs to be solved repeatedly with small perturbations in the power injections ($\mathbf{b}$), known as hot-starting of DCPF [3]. Also, the error is expressed in terms of the energy norm, $\|e^{i}\|_{A}$ (see supplementary information for more details on error definitions.). [^5] Note that, $\lambda_{1}||e||_{2}\leq\|e^{i}\|_{A}=\sqrt{(e^{i})^{T}Ae^{i}}\leq\lambda_{N}\|e^{i}\|_{2}$ iff $A$ is PSD where $\lambda_{1}<\dots<\lambda_{N}$ is the ordered set of eigenvalues of $A$. For sparse matrices, this maximum eigenvalue can be often bounded by a constant. Moreover, for multiple power systems, the susceptance matrix has eigenvalues less than one (30, 118, 200, 500, 1354, and 2383-Bus system from [35]). More details on effect of error definition on CG complexity are given with supplementary information. Therefore, the effects of the definition of $\varepsilon_{c}$ on runtime complexity must be considered carefully while comparing CG to any other classical or quantum algorithm for solving (1).
@@ IV Quantum Advantage Analysis of HHL-QPF
We now examine the complexity of different components of HHL-QPF. [^6] Although the discussion presented in this section is centered around using HHL for QPF, our analysis in general applies to other quantum linear solvers as well. Table 1 shows the complexity of another quantum algorithm to solve a linear system of equations called VTAA-HHL [14]. Complexities of state-of-the-art quantum algorithms for linear systems show linear scaling with condition number. They are also subject to the same state preparation and read-out considerations that we discuss in this work [36, 12]. The objective is to systematically analyze the complexity of the operations employed in solving QPF, culminating in an understanding of the end-to-end complexity of solving the DCPF problem.
IV.1 Reading the Problem: State Preparation Complexity
@@
To solve DCPF using a quantum computer, the problem must first be represented in terms of quantum states. Note that a vector of $N$ elements can be represented using $n=\lceil\log(N)\rceil$ qubits achieving seemingly exponential compression in memory [10, 11]. The process of transferring data from classical memory to a quantum computer is referred to as state preparation or encoding. Quantum state preparation is accomplished through the manipulation of qubits, converting classical information into a quantum superposition. The choice of encoding method depends on the particular quantum algorithm (such as HHL in the case of QPF) used to solve the problem. The encoding type is also closely tied to the performance of the particular algorithm. In the context of HHL, the technique employed for preparing a quantum state with information from vector $\mathbf{b}$ is known as amplitude encoding. In the amplitude encoding technique, data is encoded into the amplitudes of a quantum state. A normalized classical $N$ element power injection vector $\mathbf{b}$ is represented by the amplitudes of a $n$-qubit quantum state $\left|\mathbf{b}\right\rangle$ as $\left|\mathbf{b}\right\rangle=\sum_{i=1}^{N}b{}_{i}|i\rangle$ here, $b_{i}$ is the $i$-th element of normalized injection vector $b_{norm}$ (normalized injection at $i$-th bus), and $|i\rangle$ is the $i$-th computational basis state [20, 37].
@@
The QPF problem demands the ability to perform arbitrary state preparation for amplitude embedding. This necessity arises because the injeciton vector $\mathbf{b}$ lacks a specific structure, and individual injection values are represented as arbitrary floating-point numbers. Further, to prepare an arbitrary $n$-qubit state, at least $N=2^{n}$ elementary quantum gates are required, i.e., gate complexity is lower bounded by $\Omega(N)$ [12, 38]. Assuming that each unitary gate operation (one or two-qubit gate) takes unit time, the runtime complexity to prepare an arbitrary quantum state $\mathcal{O}(T_{b})$, for the injection vector is $\mathcal{O}(T_{b})=\mathcal{O}(N)$[12].
@@
An inventive method for accessing arbitrary classical data in a QC is to utilize a theorized quantum random access memory (QRAM), allowing for coherent access to classical information. The proposal to introduce and employ QRAM addresses challenges linked to state preparation or limitations in inputting data in quantum linear algebra applications [16, 12]. With the caveats on QRAM’s practicality, it can be demonstrated that the circuit depth (and runtime complexity) for accessing data from QRAM will be $\mathcal{O}(T_{b})=\mathcal{O}(\log(N))$ [12]. However, there is a one-time cost associated with initially loading the QRAM with data. Moreover, in practice, the QRAM may also cause a run-time increase due to error correction and fault tolerance overheads [39]. Further, as shown in Figure 3 (right), load vector is not sparse in nature and the majority of PGLib systems have 40-80% non-zero load values. Therefore, exploiting sparsity to encode $\mathbf{b}$ will not be beneficial.
IV.2 Solving the Problem: State Propagation Complexity
@@
The quantum mechanics equivalent of solving linear system (1) is to construct $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$. To achieve this, HHL relies on the fact that if $A$ is Hermitian, i.e., $A=A^{\dagger}$, then $A$ can be considered as a Hamiltonian for a quantum system, and the time evolution operator of such a quantum system is given by $e^{iAt}$ [13]. Other stages involved in the solution process include quantum phase estimation, ancilla bit rotation, and inverse quantum phase estimation.
@@
The quantum simulation complexity in HHL [13] is given by $\mathcal{O}(t_{o}s^{2}\log(N))$, where $t_{o}=\mathcal{O}(\kappa/\varepsilon_{h})$ is chosen to achieve the final error $\varepsilon_{h}$. Consequently, quantum simulation can be executed in $\mathcal{O}(\log(N)\kappa s^{2}/\varepsilon_{h})$ operations. In HHL, it is necessary to repeat this solution procedure to attain the ancilla qubit as $|1\rangle$ with high probability. This requires $\mathcal{O}(\kappa)$ repetitions with amplitude amplification [40] to boost the final success probability of the algorithm. Combining these terms, the overall runtime complexity of HHL, given an initial vector state $|b\rangle$, is $\mathcal{O}(\kappa^{2}s^{2}\log(N)/\varepsilon_{h})$ [13]. However, as mentioned earlier, a quantum state for the injection vector $\mathbf{b}$ needs preparation, introducing $\mathcal{O}(T_{b})$ complexity. Additionally, this state must be prepared for each $\mathcal{O}(\kappa)$ repetitions required in the amplitude amplification. This is because once a quantum state is measured, it is destroyed (a phenomenon known as wave function collapse in quantum mechanics) [10]. Therefore, the overall complexity of HHL, considering state preparation, is $\mathcal{O}(\kappa(T_{b}+\kappa s^{2}\log(N)/\varepsilon_{h}))$, where $T_{b}$ either refers to the complexity of QRAM or that of a circuit for arbitrary state preparation to encode $\mathbf{b}$ into $|\mathbf{b}\rangle$.
@@
A significant difference in HHL complexity and CG complexity is in the error definition. In CG $\varepsilon_{c}$ is defined as energy norm of initial and final errors [9] (details in supplementary information) while $\varepsilon_{h}$ is defined as the two-norm of the difference between the HHL solution state $|\tilde{\mathbf{x}}\rangle$ and the exact solution state $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$, denoted as $\varepsilon_{h}=||\tilde{\mathbf{x}}\rangle-|\mathbf{x}\rangle||_{2}$ [41]. It’s crucial to note that $\varepsilon_{h}$ does not signify the error between two classically readable solution vectors. A zero HHL error $\varepsilon_{h}$ (if achieved) would imply that HHL has ‘constructed’ the exact quantum state representing $A^{-1}|\mathbf{b}\rangle$ and not the exact classical solution.
[…]
@@
In the DCPF (and subsequently QPF) problem, our interest lies in knowing all the entries of the vector $\mathbf{x}$, representing bus angles. Therefore, the next subsection presents considerations on the complexity associated with quantum measurements or readout to obtain a classical description $\widehat{\mathbf{x}}$ of the HHL solution $|\widetilde{\mathbf{x}}\rangle$.
IV.3 Reading the Result: Measurement Complexity
@@
Upon solving a linear system using quantum computers, a classically readable vector $\widehat{\mathbf{x}}$ can only be obtained by measuring the quantum state solution $|\tilde{\mathbf{x}}\rangle$. These quantum measurements are probabilistic in nature and the probability of an outcome is governed by Born’s rule [10]. The process of obtaining a full classical description of the given quantum state is called quantum tomography [11]. To perform quantum tomography for obtaining the bus angle vector estimate $\widehat{\mathbf{x}}$, we require access to multiple copies of corresponding HHL solution $|\tilde{\mathbf{x}}\rangle$. Further, the properties of QPF dictate that we perform independent and identical measurements on each copy of HHL solution $|\tilde{\mathbf{x}}\rangle$.To obtain one copy of $|\tilde{\mathbf{x}}\rangle$, one must solve the linear system using HHL once. Therefore to obtain the end-to-end complexity of solving QPF using HHL, we need to know how many copies of $|\tilde{\mathbf{x}}\rangle$ are needed for readout to obtain a classical description of solution $\widehat{\mathbf{x}}$. Using the state-of-art results by considering that we have access to a unitary [^8] We can consider access to a unitary because $|\tilde{\mathbf{x}}\rangle$ is the output of a quantum algorithm [42]. Further, with this consideration, we require fewer copies of $|\tilde{\mathbf{x}}\rangle$, thus obtaining a lower overall complexity. (also the inverse of it) which prepares the state, we require $\Theta(\operatorname*{poly}(N/\varepsilon_{t}))$ copies of $|\tilde{\mathbf{x}}\rangle$ with $\varepsilon_{t}$ defined using $L_{2}$-norm. Here, $\Theta(\cdot)$ refers to simultaneous lower and upper bounds on asymptotic query complexity. Note that $\varepsilon_{t}\in[0,1]$ dictates the distance between a quantum state $|\tilde{\mathbf{x}}\rangle$ and its classical description $\widehat{\mathbf{x}}$. Considering that $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$ represents the exact solution, analogous to $\mathbf{x}^{\star}$ in classical CG setting, the net error in solving QPF is a combination of HHL error $\varepsilon_{q}$ and $\varepsilon_{t}$. However, discussion on the exact expression of QPF error and comparison with classical error $\varepsilon_{c}$ has been omitted considering that both CG and HHL achieve final errors within an acceptable limit and error does not scale with system size $N$.
Table 1: QPF Solving Complexity Comparison in $\mathcal{O}(\cdot)$
Algorithm Solving Complexity E2E Complexity Optimistic E2E ${}^{\#}$ with $\kappa=N^{\beta}$
CG [9] $sN\sqrt{\kappa}\log(1/\varepsilon_{c})$ $Ns+sN\sqrt{\kappa}\log(1/\varepsilon_{c})+N$ $s\sqrt{\kappa}N\log(N)\log(1/\varepsilon)$ $sN^{1+0.5\beta}\log(N)\log(1/\varepsilon)$
@@
${}^{\#}$ Considering QRAM is available and $T_{b}=\log(N)$;     E2E: End-to-End;     Discussion on $\varepsilon_{c}$ is included in supplementary material
IV.4 End-to-End Complexity
Now we analyze the end-to-end complexity of solving QPF. Starting from a state $|\mathbf{b}\rangle$, the depth of the HHL circuit required to prepare the solution $|\mathbf{\tilde{x}}\rangle$ with error $\varepsilon$ is $T_{s}=\mathcal{O}(\log(N)\kappa s^{2}/\varepsilon.)$ This circuit only has a small success probability, hence this algorithm requires amplitude amplification with $\mathcal{O}(\kappa)$ repetitions of this circuit to succeed with high probability. Let $T_{r}$ be the complexity of readout with error $\varepsilon$, assuming state-of-the-art quantum tomography methods this has a minimum complexity of $\mathcal{O}(N/\varepsilon)$ [42]. Let $T_{b}$ be the complexity of preparing the state $|\mathbf{b}\rangle$. Assuming no access to QRAM this complexity cannot be lower than $\mathcal{O}(N).$ If we have a QRAM already loaded with this state, then this complexity becomes $\mathcal{O}(\log(N)).$
Combining all these the end-to-end complexity of QPF becomes $\mathcal{O}(T_{r}(\kappa(T_{b}+T_{s})))$. This is pictorially represented in Figure 4. If we do not have QRAM access, optimistic estimates for this complexity become,

### v3

@@ III Classical Method Complexity
The conjugate gradient (CG) method is among the best algorithms in terms of asymptotic complexity for solving linear systems of equations with a sparse, positive definite matrix $A$. As the power system matrices fulfill both conditions, we focus on analyzing CG as one of the best ways to classically solve DCPF. We define $s$ as the maximum number of non-zero values in a row, i.e., sparsity, and $\kappa$ as condition number ($\kappa=\|A^{-1}\|\|A\|$). The total complexity of CG can be split into two parts: complexity of dominant operations and number of times dominant operations are used.
Considering the dominant operation complexity of matrix-vector multiplication, $\mathcal{O}(Ns),$ and an upper bound on number of iterations required to achieve the error $\varepsilon_{c}$ (details are given in appendix), the overall complexity of the CG method for solving (1) is $\mathcal{O}(N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}}))$. Further, read-write complexity for the sparse matrix $A$ is $\mathcal{O}(Ns)$ (higher than the $\mathbf{b},\mathbf{x}$ readout cost, which is $\mathcal{O}(N)$), and this additional cost is additive to the runtime complexity. Here, additive means that we read $A,\mathbf{b}$ once before solving (1) and read $\mathbf{x}\in\mathbb{R}^{N}$ once after solving the problem. To obtain $\mathbf{x}$, CG does not need to be repeated but only needs to be read with complexity $\mathcal{O}(N)$. This is important as additive terms are subsumed into dominant complexity terms and do not alter the end-to-end complexity. Thus, the end-to-end complexity of solving DCPF (1) using CG is $\mathcal{O}(Ns+N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}})+N)=\mathcal{O}(N\sqrt{\kappa}s\log({1}/{\varepsilon_{c}}))$. [^4] For basic Big-O notation introduction, readers can refer to these notes: Link. Along with exploiting sparsity, CG runtime complexity has a square root dependence condition number of $\kappa$, making it highly efficient for DCPF as $\kappa$ has polynomial growth, as shown in Figure 3 (center).
[…]
@@ IV Quantum Advantage Analysis of HHL-QPF
We now examine the complexity of different components of HHL-QPF. [^6] Although the discussion presented in this section is centered around using HHL for QPF, our analysis in general applies to other quantum linear solvers as well. Table I shows the complexity of another quantum algorithm to solve a linear system of equations called VTAA-HHL [14]. Complexities of state-of-the-art quantum algorithms for linear systems show linear scaling with condition number. They are also subject to the same state preparation and read-out considerations that we discuss in this work [41, 12]. The objective is to systematically analyze the complexity of the operations employed in solving QPF, culminating in an understanding of the end-to-end complexity of solving the DCPF problem.
IV-A Reading the Problem: State Preparation Complexity
@@
To solve DCPF using a quantum computer, the problem must first be represented in terms of quantum states. A vector of $N$ elements can be represented using $n=\lceil\log(N)\rceil$ qubits, achieving seemingly exponential compression in memory [10, 11]. The process of transferring data from classical memory to a quantum computer is referred to as state preparation or encoding. Quantum state preparation is accomplished through the manipulation of qubits, converting classical information into a quantum superposition. The choice of encoding method depends on the particular quantum algorithm (such as HHL in the case of QPF) used to solve the problem. The encoding type is also closely tied to the performance of the particular algorithm. In the context of HHL, the technique employed for preparing a quantum state with information from vector $\mathbf{b}$ is known as amplitude encoding. In the amplitude encoding technique, data is encoded into the amplitudes of a quantum state. A normalized classical $N$ element power injection vector $\mathbf{b}$ is represented by the amplitudes of an $n$-qubit quantum state $\left|\mathbf{b}\right\rangle$ as $\left|\mathbf{b}\right\rangle=\sum_{i=1}^{N}b_{i}|i\rangle$. Here, $b_{i}$ is the $i$-th element of normalized injection vector $b_{norm}$ (normalized injection at $i$-th bus), and $|i\rangle$ is the $i$-th computational basis state [21, 37].
@@
The QPF problem demands the ability to perform arbitrary state preparation for amplitude embedding. This necessity arises because the injection vector $\mathbf{b}$ lacks a specific structure, and individual injection values are represented as arbitrary floating-point numbers. Further, to prepare an arbitrary $n$-qubit state, at least $N=2^{n}$ elementary quantum gates are required, i.e., gate complexity is lower bounded by $\Omega(N)$ [12, 42]. Assuming that each unitary gate operation (one or two-qubit gate) takes unit time, the runtime complexity to prepare an arbitrary quantum state $\mathcal{O}(T_{b})$ for the injection vector is $\mathcal{O}(T_{b})=\mathcal{O}(N)$[12].
@@
An inventive method for accessing arbitrary classical data in a QC is to use a theorized quantum random access memory (QRAM), allowing for coherent access to classical information. The proposal to introduce and employ QRAM addresses challenges linked to state preparation or limitations in inputting data in quantum linear algebra applications [16, 12]. With those caveats on QRAM’s practicality, the circuit depth (and runtime complexity) for accessing data from QRAM is demonstrated to be $\mathcal{O}(T_{b})=\mathcal{O}(\log(N))$ [12]. However, there is a one-time cost associated with initially loading QRAM with data. Moreover, in practice, QRAM may also cause a runtime increase due to error correction and fault tolerance overheads [43]. Further, as shown in Figure 3 (right), the load vector is not sparse in nature and the majority of PGLib systems have 40-80% non-zero load values. Therefore, exploiting sparsity to encode $\mathbf{b}$ will not be beneficial.
IV-B Solving the Problem: State Propagation Complexity
@@
The quantum mechanics equivalent of solving linear system (1) is to construct $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$. To achieve this, HHL relies on the fact that if $A$ is Hermitian, i.e., $A=A^{\dagger}$, then $A$ can be considered a Hamiltonian for a quantum system, and the time evolution operator of such a quantum system is given by $e^{iAt}$ [13]. Other stages involved in the solution process include quantum phase estimation, ancilla bit rotation, and inverse quantum phase estimation.
@@
The quantum simulation complexity in HHL [13] is given by $\mathcal{O}(t_{o}s^{2}\log(N))$, where $t_{o}=\mathcal{O}(\kappa/\varepsilon_{h})$ is chosen to achieve the final error $\varepsilon_{h}$. Consequently, quantum simulation can be executed in $\mathcal{O}(\log(N)\kappa s^{2}/\varepsilon_{h})$ time. In HHL, it is necessary to repeat this solution procedure to attain the ancilla qubit as $|1\rangle$ with high probability. This requires $\mathcal{O}(\kappa)$ repetitions with amplitude amplification [44] to boost the final success probability of the algorithm. Combining these terms, the overall runtime complexity of HHL, given an initial vector state $|b\rangle$, is $\mathcal{O}(\kappa^{2}s^{2}\log(N)/\varepsilon_{h})$ [13]. However, as mentioned earlier, a quantum state for the injection vector $\mathbf{b}$ needs preparation, introducing $\mathcal{O}(T_{b})$ complexity. Additionally, this state must be prepared for each $\mathcal{O}(\kappa)$ repetition required in the amplitude amplification. This is because once a quantum state is measured, it is destroyed (a phenomenon known as wave function collapse in quantum mechanics) [10]. Therefore, the overall complexity of HHL, considering state preparation, is $\mathcal{O}(\kappa(T_{b}+\kappa s^{2}\log(N)/\varepsilon_{h}))$, where $T_{b}$ either refers to the complexity of QRAM or that of a circuit for arbitrary state preparation to encode $\mathbf{b}$ into $|\mathbf{b}\rangle$.
@@
A significant difference between HHL complexity and CG complexity is in error definition. In CG, $\varepsilon_{c}$ is defined as the energy norm of initial and final errors [9] (details in appendix) while $\varepsilon_{h}$ is defined as the two-norm of the difference between the HHL solution state $|\tilde{\mathbf{x}}\rangle$ and the exact solution state $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$, denoted as $\varepsilon_{h}=||\tilde{\mathbf{x}}\rangle-|\mathbf{x}\rangle||_{2}$ [45]. It’s crucial to note that $\varepsilon_{h}$ does not signify the error between two classically readable solution vectors. A zero HHL error $\varepsilon_{h}$ (if achieved) would imply that HHL has ‘constructed’ the exact quantum state representing $A^{-1}|\mathbf{b}\rangle$ and not the exact classical solution.
If we analyze the runtime complexity of HHL without considering $T_{b}$, we observe that HHL exhibits exponential speedup with respect to the system size $N$, quadratic slowdown in sparsity $s$, and the condition number $\kappa$ [45] in comparison to CG. Advancements in quantum linear system-solving methods have further widened this perceived quantum speedup from CG [12]. However, these runtime complexities do not capture the time required to obtain a classically readable and usable solution vector $\mathbf{x}$. The HHL algorithm is only suitable for problems where the user is interested in measuring a few values of the form $\langle\mathbf{x}|M|\mathbf{x}\rangle$ [^7] MM is a linear operator and HHL considers that we are interested in the expectation value of $M$ acting on $\widehat{x}$, the solution vector. from the solution state $|\mathbf{x}\rangle$.
@@
In the DCPF (and subsequently QPF) problem, our interest lies in knowing all the entries of the vector, $\mathbf{x}$, representing bus angles. Therefore, the next subsection presents considerations of the complexity associated with quantum measurements or readout to obtain a classical description $\widehat{\mathbf{x}}$ of the HHL solution $|\widetilde{\mathbf{x}}\rangle$.
IV-C Reading the Result: Measurement Complexity
@@
Upon solving a linear system using quantum computers, a classically readable vector $\widehat{\mathbf{x}}$ can only be obtained by measuring the quantum state solution $|\tilde{\mathbf{x}}\rangle$. These quantum measurements are probabilistic in nature and the probability of an outcome is governed by Born’s rule [10]. The process of obtaining a full classical description of the given quantum state is called quantum tomography [11]. To perform quantum tomography for obtaining the bus angle vector estimate $\widehat{\mathbf{x}}$, access to multiple copies of corresponding HHL solution $|\tilde{\mathbf{x}}\rangle$ is required. Further, the properties of QPF dictate that independent and identical measurements be performed on each copy of an HHL solution $|\tilde{\mathbf{x}}\rangle$.To obtain one copy of $|\tilde{\mathbf{x}}\rangle$, linear system must be solved using HHL once. Therefore, to obtain the end-to-end complexity of solving QPF using HHL, we need to know how many copies of $|\tilde{\mathbf{x}}\rangle$ are needed for readout to obtain a classical description of the solution $\widehat{\mathbf{x}}$. Using state-of-art results by considering that we have access to a unitary [^8] We can consider access to a unitary because $|\tilde{\mathbf{x}}\rangle$ is the output of a quantum algorithm [46]. Further, with this consideration, we require fewer copies of $|\tilde{\mathbf{x}}\rangle$, thus obtaining a lower overall complexity. (also the inverse of it) that prepares the state, we require $\Theta(\operatorname*{poly}(N/\varepsilon_{t}))$ copies of $|\tilde{\mathbf{x}}\rangle$ with $\varepsilon_{t}$ defined using $L_{2}$-norm. Here, $\Theta(\cdot)$ refers to simultaneous lower and upper bounds on asymptotic query complexity. Note that $\varepsilon_{t}\in[0,1]$ dictates the distance between a quantum state $|\tilde{\mathbf{x}}\rangle$ and its classical description $\widehat{\mathbf{x}}$, obtained through multiple readouts. Considering that $|\mathbf{x}\rangle=A^{-1}|\mathbf{b}\rangle$ represents the exact solution and is analogous to $\mathbf{x}^{\star}$ in the classical CG setting, the net error in solving QPF is a combination of HHL errors $\varepsilon_{q}$ and $\varepsilon_{t}$. Discussion on the exact expression of QPF error and comparison with classical error $\varepsilon_{c}$ is omitted considering that both CG and HHL achieve final errors within an acceptable limit and error does not scale with system size $N$.
TABLE I: QPF Solving Complexity Comparison in $\mathcal{O}(\cdot)$
Algorithm Solving Complexity E2E Complexity Optimistic E2E Complexity# with $\kappa=N^{\beta}$
CG [9] $\sqrt{\kappa}\log(1/\varepsilon_{c}) Ns+sN$$sN\sqrt{\kappa}\log(1/\varepsilon_{c})+N$ $s\sqrt{\kappa}N\log(N)\log(1/\varepsilon)$ $sN^{1+0.5\beta}\log(N)\log(1/\varepsilon)$
@@
Lowerbound [13, 47] $\kappa\log(1/\varepsilon)$ $\kappa N\log(1/\varepsilon)/\varepsilon$ $N^{1+\beta}(\log(1/\varepsilon)/\varepsilon)$
# Considering QRAM is available and $T_{b}=\log(N)$;   E2E: End-to-End;   Discussion on $\varepsilon_{c}$ is included in appendix
IV-D End-to-End Complexity
Now we analyze the end-to-end complexity of solving QPF. Starting from a state $|\mathbf{b}\rangle$, the depth of the HHL circuit required to prepare the solution $|\mathbf{\tilde{x}}\rangle$ with error $\varepsilon$ is $T_{s}=\mathcal{O}(\log(N)\kappa s^{2}/\varepsilon)$. This circuit only has a small success probability, therefore, the algorithm requires amplitude amplification with $\mathcal{O}(\kappa)$ repetitions of this circuit to succeed with high probability. Let $T_{r}$ be the complexity of readout with error $\varepsilon$. Assuming state-of-the-art quantum tomography methods, this complexity has a minimum complexity of $\mathcal{O}(N/\varepsilon)$ [46]. Let $T_{b}$ be the complexity of preparing the state $|\mathbf{b}\rangle$. Assuming no access to QRAM, this complexity cannot be lower than $\mathcal{O}(N).$ If we have a QRAM already loaded with this state, then this complexity becomes $\mathcal{O}(\log(N)).$
Combining all of these complexities, the end-to-end complexity of QPF becomes $\mathcal{O}(T_{r}(\kappa(T_{b}+T_{s})))$. This is pictorially represented in Figure 4. If we do not have QRAM access, optimistic estimates for this complexity become